# ToolPlanner: A Tool Augmented LLM for Multi Granularity Instructions with Path Planning and Feedback

### v1

@@ Abstract
Recently, tool-augmented LLMs have gained increasing attention. Given an instruction, tool-augmented LLMs can interact with various external tools in multiple rounds and provide a final answer. However, previous LLMs were trained on overly detailed instructions, which included API names or parameters, while real users would not explicitly mention these API details. This leads to a gap between trained LLMs and real-world scenarios. In addition, most works ignore whether the interaction process follows the instruction. To address these issues, we constructed a training dataset called MGToolBench, which contains statement and category-level instructions to better reflect real-world scenarios. In addition, we propose ToolPlanner, a two-stage reinforcement learning framework that utilizes path planning and two feedback mechanisms to enhance the LLM’s task completion and instruction-following capabilities. Experimental results show that ToolPlanner significantly improves the Match Rate, Pass Rate and Win Rate by 26.8%, 20.2%, and 5.6% compared to the SOTA model. Human evaluation verifies that the multi-granularity instructions can better align with users’ usage habits. Our data and code will be released upon acceptance.
@@ 4.4 Training
ChatGPT-Chain 4 22 17 13 11 7 21 15 8 13.1 42 25 36 31 20 30.8 50 28 44 34 32 37.6
ChatGPT-Tree 3 20 14 15 14 8 16 11 4 11.7 55 32 46 39 38 42.0 57 43 48 46 45 47.8
GPT4-Chain 8 51 42 25 23 18 39 34 21 29.0 71 60 68 62 41 60.4 82 80 71 74 54 72.2
GPT4-Tree 10 44 38 23 21 14 45 40 24 28.8 67 63 68 64 45 61.4 74 79 66 76 63 71.6
ToolLLaMA-Chain 5 48 35 30 24 13 41 34 9 26.6 61 30 41 20 14 33.2 69 41 43 28 47 45.6
ToolPlanner 59 61 57 64 61 52 60 52 37 55.8 88 89 84 83 78 84.4 78 79 77 80 75 77.8
@@
5 Experiment
5.1 Dataset
5.2 Settings
5.3 Results and Discussions
The main experimental results for Match Rate, Pass Rate, and Win Rate are presented in Table 1. From the table we can observe that:
$\bullet$ Overall, ToolPlanner improves the Match Rate, Pass Rate, and Win Rate by +26.8%, +20.2% and +5.6%. This indicates that ToolPlanner can flexibly complete instructions of different levels, and provide high-quality answers for instructions that are close to real-world scenarios.
Level Plausibility Conciseness Relevance Realness
Statement 2.85 2.99 - 2.94
Category 2.73 2.97 2.60 2.89
Tool 2.58 2.98 2.50 2.68
API 2.01 2.72 2.13 2.19
Hybrid 2.62 2.01 2.46 2.60
@@ 5.4 Human Evaluation
To evaluate our multi-granularity instruction mechanism, we conducted a human evaluation. We randomly selected 100 instructions with different levels and asked three annotators to rate them on a 1-3 scale across four metrics. As shown in Table 3:
@@ B.3.3 Prompt Design
The prompts of the category-level, tool-level and API-level instruction generation are shown in Table 13, 14, 15, respectively.
@@ C.1.1 Prompt Design
The prompts of tag extraction, solution path planning and solution tree generation are shown in Table 16, 17, 18, respectively.

### v2

@@ Abstract
Recently, tool-augmented LLMs have gained increasing attention. Given an instruction, tool-augmented LLMs can interact with various external tools in multiple rounds and provide a final answer. However, previous LLMs were trained on overly detailed instructions, which included API names or parameters, while real users would not explicitly mention these API details. This leads to a gap between trained LLMs and real-world scenarios. In addition, most works ignore whether the interaction process follows the instruction. To address these issues, we constructed a training dataset called MGToolBench, which contains statement and category-level instructions to better reflect real-world scenarios. In addition, we propose ToolPlanner, a two-stage reinforcement learning framework that utilizes path planning and two feedback mechanisms to enhance the LLM’s task completion and instruction-following capabilities. Experimental results show that ToolPlanner significantly improves the Match Rate, Pass Rate and Win Rate by 26.8%, 20.2%, and 5.6% compared to the SOTA model. Human evaluation verifies that the multi-granularity instructions can better align with users’ usage habits. Our data and code are available at https://github.com/XiaoMi/toolplanner.
@@ 4.4 Training
ChatGPT-Chain FewShot 4 22 17 13 11 7 21 15 8 13.1 42 25 36 31 20 30.8 50 28 44 34 32 37.6
ChatGPT-Tree FewShot 3 20 14 15 14 8 16 11 4 11.7 55 32 46 39 38 42.0 57 43 48 46 45 47.8
GPT4-Chain FewShot 8 51 42 25 23 18 39 34 21 29.0 71 60 68 62 41 60.4 82 80 71 74 54 72.2
GPT4-Tree FewShot 10 44 38 23 21 14 45 40 24 28.8 67 63 68 64 45 61.4 74 79 66 76 63 71.6
ToolLLaMA-Chain SFT 5 48 35 30 24 13 41 34 9 26.6 61 30 41 20 14 33.2 69 41 43 28 47 45.6
ToolLLaMA-Tree SFT 8 33 22 26 19 9 38 28 13 21.8 68 60 74 66 53 64.2 73 64 77 73 69 71.2
ToolLLaMA-2-7b SFT 10 47 41 33 29 16 50 45 26 33 82 74 85 74 60 75 79 76 77 73 69 74.8
ToolPlanner SFT 59 61 57 64 61 52 60 52 37 55.8 88 89 84 83 78 84.4 78 79 77 80 75 77.8
@@
Model Method Pass Win
GPT-4 Turbo Chain Guo et al. (2024) FewShot 52.5 67.2
GPT-4 Turbo Tree Guo et al. (2024) FewShot 66.1 60.7
ToolLLaMA v2 Chain Guo et al. (2024) SFT 33.9 24.6
ToolLLaMA v2 Tree Guo et al. (2024) SFT 53.6 50.8
AnyTool Du et al. (2024) Agent 63.2 -
Sum2Act Liu et al. (2024) Agent 74 50.6
ToolPlanner SFT 78 75
Table 2: Pass Rate and Win Rate of previous studies.
5 Experiment
5.1 Dataset
5.2 Settings
5.3 Results and Discussions
The main experimental results for Match Rate, Pass Rate, and Win Rate are presented in Table 1 and Table 2. From the table we can observe that:
$\bullet$ Overall, ToolPlanner improves the Match Rate, Pass Rate, and Win Rate by +26.8%, +20.2% and +5.6% compared to ToolLLaMA-Tree. This indicates that ToolPlanner can flexibly complete instructions of different levels, and provide high-quality answers for instructions that are close to real-world scenarios.
Level Plausibility Conciseness Relevance Realness
Statement 2.85 2.99 - 2.94
Category 2.73 2.97 2.60 2.89
Tool 2.58 2.98 2.50 2.68
API 2.01 2.72 2.13 2.19
Hybrid 2.62 2.01 2.46 2.60
Model Method Human Human Statement
Match Pass Win Match Pass Win
Category Tool API Category Tool API
GPT4 FewShot 45 39 9 70 78 42 36 8 55 76
ToolLLaMA ToolBench SFT 46 36 8 69 75 40 34 10 56 73
ToolLLaMA-v2-7b ToolBench SFT 59 54 13 74 79 57 48 11 71 75
ToolLLaMA-SFT MGToolBench SFT w/o RL 25 25 11 29 31 31 27 9 36 30
ToolLLaMA-2-SFT MGToolBench SFT w/o RL 49 46 15 70 64 40 34 10 66 63
ToolPlanner MGToolBench SFT 60 54 20 75 77 55 50 17 84 75
Table 4: Model performance on 100 human-written instructions.
@@ 5.4 Human Evaluation
Quality of multi-granularity instructions: To evaluate our multi-granularity instruction mechanism, we conducted a human evaluation. We randomly selected 100 instructions with different levels and asked three annotators to rate them on a 1-3 scale across four metrics. As shown in Table 3:
@@
Performance on human-written instructions: To verify the effectiveness of the model on real-world user instructions, we ask human annotators to handwrite new instructions based on the tag lists from the test set as a real-world test set. We believe that these human-written instructions can better reflect real-world scenarios. Specifically, there are two types of human-written instructions:
1. Human: annotators write instructions based only on the tag lists.
2. Human-Statement: annotators can use both statement sentences and tag lists.
We use different fine-tuning baselines to generate solutions based on these human-written instructions. The evaluation results are shown in Table 4. These results show that our ToolPlanner still outperforms the rest baselines on Match Rate and Pass Rate, and is slightly worse than GPT-4 on Win Rate. We believe this proves the effectiveness of ToolPlanner in real-world user instructions.
@@ B.3.3 Prompt Design
The prompts of the category-level, tool-level and API-level instruction generation are shown in Table 15, 16, 17, respectively.
@@ C.1.1 Prompt Design
The prompts of tag extraction, solution path planning and solution tree generation are shown in Table 18, 19, 20, respectively.